OpenStudy (anonymous):

If the point (3\5, 4\5) corresponds to angle Θ in the unit circle, what is tan Θ? 4\3 3\4 5\4 5\3

OpenStudy (anonymous):

its 5/4

OpenStudy (anonymous):

the y co-ordinate is 4/5, x-co-ordinate is 3/5, so tan theta=y/x=4/3

OpenStudy (anonymous):

Thank you!